OpenStudy (anonymous):

Three forces in the x-y pkane act on a 5.70Kg mass. 16.90N directed at 12deg, 13.40N directed at 116deg, and 5.90N directed at 243deg. Calculate de magnitud of the acceleration. Help!!!!!!

OpenStudy (michele_laino):

please I rewrite all of your three forces using their components: \[\begin{gathered} {{\mathbf{F}}_{\mathbf{1}}} = \left( {16.9\cos \left( {12} \right),16.9\sin \left( {12} \right)} \right) \hfill \\ {{\mathbf{F}}_{\mathbf{2}}} = \left( {13.4\cos \left( {116} \right),13.4\sin \left( {116} \right)} \right) \hfill \\ {{\mathbf{F}}_{\mathbf{3}}} = \left( {5.9\cos \left( {243} \right),5.9\sin \left( {243} \right)} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, their resultant is: \[{{\mathbf{F}}_{\mathbf{1}}} + {{\mathbf{F}}_{\mathbf{2}}} + {{\mathbf{F}}_{\mathbf{3}}}\] and the x-component of that resultant forcre is: \[16.9\cos \left( {12} \right) + 13.4\cos \left( {116} \right) + 5.9\cos \left( {243} \right)\] whereas the y-component of that resultant is: \[16.9\sin \left( {12} \right) + 13.4\sin \left( {116} \right) + 5.9\sin \left( {243} \right)\]